A rocket lifts off from ground and accelerate upwards at 1ms^(2).20 seconds after lift off a piece breaks off from the bottom of rocket.After breaking off how much time it takes approximately to reach the ground? (Take g=10ms^(2) )
Ans 8.5 s
1
Expert's answer
2021-08-30T15:06:17-0400

"\u03c5 = \u03c5_0+ at"

"y = y_0 + v_0t+\frac{at^2}{2}"

"\text{For a rocket through }t=20s"

"a = 1\frac{m}{s^2}"

"v_0 = 0"

"v_r = 1*20= 20\frac{m}{s}"

"y_0 = 0"

"y_r = \frac{1*20^2}{2}=200m"

"\text{For a piece of rocket}"

"a =- g = -10\frac{m}{s^2}"

"v_0= v_r = 20\frac{m}{s}"

"y_p = 0"

"y_0 = y_r = 200m"

"y = y_0 + v_0t+\frac{at^2}{2}"

"0 = 200 +20t-\frac{10*t^2}{2}"

"t^2- 4t-40 = 0"

"t_1 \approx 8.63 s"

"t_2 \approx -4.63 <0"


"\text{Answer: }8.63s"